OpenStudy (anonymous):

Find x and y in 3x + yi = 5x + 1 + 2i

OpenStudy (anonymous):

here u hqve to equate real parts of one side with the real part of the other side and same for imaginary part on both side hence we have 3x=5x+1 and y=2 or x=-1/2 and y=2

OpenStudy (anonymous):

:D Good job

OpenStudy (anonymous):

Time to close :P